In which of the following is y a function of x
a. 3y+2x-9=0
b. x+y^2=5
c. x^2+y^2=25
d. noone of these

Answer by venugopalramana(3286)

In which of the following is y a function of x
a. 3y+2x-9=0.....3Y=9-2X.....Y=(9-2X)/3..WE GET UNIQUE VALUE FOR Y .SO IT IS A FUNCTION
b. x+y^2=5.....Y^2=5-X....Y=+ OR - SQRT(5-X)..WE DO NOT GET UNIQUE VALUE FOR Y .SO IT IS NOT A FUNCTION
c. x^2+y^2=25....Y^2=25-X^2....Y=+ OR - SQRT(25-X^2)..WE DO NOT GET UNIQUE VALUE FOR Y .SO IT IS NOT A FUNCTION
d. noone of these....NOT CORRECT....A..IS THE ANSWER